1. Consider a call option selling for $ 4 in which the exercise price is $50.
A) Determine the value at expiration and the profit for a buyer under the following outcomes:
i. The price of the underlying at expiration is $55
ii. The price of the underlying at expiration is $51
iii. The price of the underlying at expiration is $48
B) Determine the value at expiration and the profit for a seller under the following outcomes:
i. The price of the underlying at expiration is $49
ii. The price of the underlying at expiration is $52
iii. The price of the underlying at expiration is $55
C) Determine the following:
i. The max profit to the buyer (max loss to the seller)
ii. The max loss to the seller (max profit to seller)
D) Determine the breakeven price of the underlying at expiration
2. Consider a put option on the Nasdaq 100 selling for $106.25 in which exercise price is 2100.
A) Determine the value at expiration and the profit for a buyer under the following outcomes:
i. The price of the underlying at expiration is 2125
ii. The price of the underlying at expiration is 2050
iii. The price of the underlying at expiration is 1950
B) Determine the value at expiration and the profit for a seller under the following outcomes:
i. The price of the underlying at expiration is 1975
ii. The price of the underlying at expiration is 2150
C) Determine the following:
iii. The max profit to the buyer (max loss to the seller)
iv. The max loss to the seller (max profit to seller)
D) Determine the breakeven price of the underlying at expiration
3. You simultaneously purchase an underlying priced at $77 and write a call option on it with an exercise price of $80 and selling at $6.
A) What is the term commonly used for the position that you have taken?
B) Determine the value at expiration and the profit for your strategy under the following outcomes:
i. The price of the underlying at expiration is $70
ii. The price of the underlying at expiration is $75
iii. The price of the underlying at expiration is $80
iv. The price of the underlying at expiration is $85
C) Determine the following:
i. The max profit
ii. The max loss
iii. The expiration price of the underlying at which you would realize the max profit
iv. The expiration price of the underlying at which you would realize the max loss
D) Determine the breakeven price at expiration
4. Suppose you simultaneously purchase an underlying priced at $77 and a put option on it with an exercise price of $75 and selling at $3.
A) What is the term commonly used for the position that you have taken?
B) Determine the value at expiration and the profit for your strategy under the following outcomes:
i. The price of the underlying at expiration is $70
ii. The price of the underlying at expiration is $75
iii. The price of the underlying at expiration is $80
iv. The price of the underlying at expiration is $85
C) Determine the following:
i. The max profit
ii. The max loss
iii. The expiration price of the underlying at which you would realize the max loss
D) Determine the breakeven price at expiration
5. You are bullish about the underlying that is currently trading at a price of $80. You choose to go long one call option on the underlying with an exercise price of $75 and selling at $10, and go short 1 call option on the underlying with an exercise price of $85 and selling at $2. Both the calls expire in 3 months.
A) What is the term commonly used for the position that you have taken?
B) Determine the value at expiration and the profit for your strategy under the following outcomes:
i. The price of the underlying at expiration is $89
ii. The price of the underlying at expiration is $78
iii. The price of the underlying at expiration is $70
C) Determine the following:
i. The max profit
ii. The max loss
D) Determine the breakeven underlying price at expiration of the call options
E) Verify that your answer to Part D above is correct
6. You expect a currency to depreciate with respect to USD. The currency is currently trading at a price of $0.75. You decide to go long 1 put option on the currency with an exercise price of $0.85 and selling at $0.15, and go short 1 put option on the currency with an exercise price of $0.70 and selling at $0.30. Both the puts expire in 3 months.
A) What is the term commonly used for the position you have taken?
B) Determine the value at expiration and the profit for your strategy under the following outcomes:
i. The price of the underlying at expiration is $0.87
ii. The price of the underlying at expiration is $0.78
iii. The price of the underlying at expiration is $0.68
C) Determine the following:
i. The max profit
ii. The max loss
D) Determine the breakeven underlying price at the expiration of the put options
E) Verify that youк answer to Part В above is correct
7. A stock is currently trading at a price of $114. You construct a butterfly spread using calls of 3 different strike prices on this stock, with the calls expiring at the same time. You go long 1 call with an exercise price of $110 and selling at $8, go short 2 calls with an exercise price of $115 and selling at $5, and go long 1 call with an exercise price of $120 and selling at $3.
A) Determine the value at expiration and the profit for your strategy under the following outcomes:
i. The price of the underlying at expiration of the calls is $106
ii. The price of the underlying at expiration of the calls is $110
iii. The price of the underlying at expiration of the calls is $115
iv. The price of the underlying at expiration of the calls is $120
v. The price of the underlying at expiration of the calls is $123
B) Determine the following:
i. The max profit
ii. The max loss
iii. The stock price at which you would realize the max profit
iv. The stock price at which you would incur the max loss
C) Determine the breakeven underlying price at expiration of the call options
Answer #1

You have asked so many questions. Each question has so many sub parts. I have answered all the the sub parts of the first question. Please post the balance questions separately.

1. Consider a call option selling for C = $ 4 in which the exercise price is K = $50.
A) Determine the value at expiration and the profit for a buyer under the following outcomes:

Value of buyer at expiration = max (S - K, 0); Profit = max (S - K, 0) - C

i. The price of the underlying at expiration is S = $55: Value = max (55 - 50, 0) = 5; Profit = 5 - 4 = 1
ii. The price of the underlying at expiration is $51; Value = 1; Profit = -3
iii. The price of the underlying at expiration is $48; Value = 0; Profit = -4
B) Determine the value at expiration and the profit for a seller under the following outcomes:

Value to seller at expiration = - max (S - K, 0); Profit = - max (S - K, 0) + C
i. The price of the underlying at expiration is $49: Value = - max (49 - 50, 0) = 0; Profit = 0 + 4 = 4
ii. The price of the underlying at expiration is $52: Value = - max (52 - 50, 0) = -2; Profit = -2 + 4 = 2
iii. The price of the underlying at expiration is $55: Value = - 5: profit = -1

C) Determine the following:
i. The max profit to the buyer (max loss to the seller) : Infinite, unlimited
ii. The max loss to the buyer (max profit to seller): $ 4

D) Determine the break even price of the underlying at expiration

Break even price = K + C = 50 + 4 = $ 54